OpenStudy (anonymous):

Find the nth term of the geometric sequence with first term -2 and common ratio 1/3.

OpenStudy (cwrw238):

nth term = a r^(n-1) where a=first term and r = common ratio

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

what is a? @cwrw238

OpenStudy (cwrw238):

first term = -2

OpenStudy (anonymous):

so a1?

OpenStudy (anonymous):

I have -2 * 1/3^(n-1)

OpenStudy (anonymous):

but what is n?

OpenStudy (cwrw238):

yes first term is sometimes referred to as first term

OpenStudy (cwrw238):

as a1!

OpenStudy (anonymous):

what is n in my equation above

OpenStudy (cwrw238):

you asked for the nth term - n is any term

OpenStudy (anonymous):

So what do I plug in there??

OpenStudy (anonymous):

I have -2 * 1/3^(n-1) this is what i have, how to solve?

OpenStudy (cwrw238):

your answer is correct -2*(1/3)^(n-1)

OpenStudy (cwrw238):

thats it - thats the formula for the nth term

OpenStudy (anonymous):

are u sure that is all I put?

OpenStudy (cwrw238):

for example if you need the sec ond term it is -2*(1/3)^(2-1) replacing n by 2

OpenStudy (cwrw238):

= -2 * 1/3 = -2/3

OpenStudy (cwrw238):

yes - i'm sure

OpenStudy (anonymous):

so my final answer is -2*(1/3)^(n-1) ?

OpenStudy (cwrw238):

right